# An Informative Planning Framework for Target Tracking and Active Mapping in Dynamic Environments with ASVs

### v2

@@ IV-D Spatiotemporal Prediction Network
We introduce a spatiotemporal prediction network to predict the uncertainty of the target positions in the map over some time horizon. Although our dynamic occupancy grid mapping method contains a prediction step to estimate the drift of targets, it is not accurate at predicting drift for longer time horizons, due to the stochastic nature of environmental disturbances. Therefore, we first create a formulation to estimate the spatial and temporal variations of the target positions for future time steps.
We assume that the position of each target is represented by a Gaussian kernel, with the uncertainty in its position estimate growing with time, similar to previous work [23]. To reflect this assumption, we assume that the variances of the kernel are proportional to the wind speed and increase with the prediction time, with the principal axis being along the direction of the measured wind. We assume that the centers of each kernel translate by distances calculated as per (6). The variance along the direction of the wind is assumed to be greater, as for most realistic scenarios, the wind speed varies faster than the direction [28]. Based on this assumption, the standard deviations of the kernel along the major and minor axes, $\sigma_{\parallel}$, and $\sigma_{\perp}$ for a prediction interval $t$ are set as $\sigma_{\parallel}=0.5\,\gamma v_{w}t$, and $\sigma_{\perp}=0.2\,\gamma v_{w}t$ respectively. In case of negligible wind speeds ($v_{w}\approx 0$), $\sigma_{\parallel}=\sigma_{\perp}=0.1\,t$.
@@
We provide the wind parameters $v_{w}$, $\psi_{w}$, and the prediction interval as other inputs to the network. The network output is a grid of the same size with predicted position uncertainties of the targets at an arbitrary time step. We use a UNet architecture [29], which is widely used for image-to-image prediction tasks such as segmentation. The wind parameters and the prediction time are embedded into the latent space of the network.
To train our network, a synthetic dataset is created to simulate input conditions with different wind conditions, target positions, and prediction intervals. To construct inputs resembling target occupancies in the map, binary grids of size $100\times 100$ containing an arbitrary number of objects (<20<20) with random positions are generated. For each of these input grids representing spatial distribution of the targets, outputs are computed as described above, for various combinations of wind parameters and prediction time steps. While predicting the drift of objects on the water surface is a difficult problem [27], we find that these assumptions suit our problem setup. Using an input representation that is agnostic to the number of targets makes our method generalizable, as validated through both simulation and field experiments, in Secs. V and VI.
@@ IV-E Informative Planning for Dynamic Target Tracking
As the environment is dynamic, the robot needs to replan frequently to maintain an accurate map of target positions. The robot must explore unknown regions of the map while redetecting targets. Planning informative trajectories for the robot involves predicting both spatial and temporal variations in the environment. Commonly used IPP objectives are based on Shannon entropy measured based on the map state [10, 25]. However, these formulations do not explicitly account for temporal variations and periodic target redetections. In contrast, we exploit the predictions from our spatiotemporal network to include the temporal uncertainties in target positions in the planning objective.
To trade-off between exploring the environment and redetecting targets, we introduce a new planning utility that consists of both exploration and target tracking objectives. We propose a new target tracking component in the utility that explicitly considers possible target detections by estimating the spatiotemporal variations in the targets’ position estimates in the planning objective. It is computed with the predictions from our spatiotemporal network, for prediction steps over the considered planning horizon. To compute the predicted change in entropy, the dynamic occupancy grid mapping is forward-simulated for predicted measurements along considered trajectories. We define the planning utility as:
$I(\mathcal{T})$ $=\Big{[}H(\mathcal{M}\,|\,z_{\mathcal{T}})-H(\mathcal{M})\Big{]}+\frac{w}{|\mathcal{T}|}\sum_{t}{J(\boldsymbol{x}^{t})}\,$, (9)
@@
where $H(\mathcal{M})$ denotes the mean Shannon entropy of the occupancy grid, and $H(\mathcal{M}|z_{\mathcal{T}})$ denotes the predicted posterior mean entropy after simulating measurements $z_{\mathcal{T}}$ along a trajectory $\mathcal{T}$. The target tracking objective is defined as the averaged sum of the utility $J(\boldsymbol{x}^{t})$ over the prediction steps, where $\boldsymbol{x}^{t}$ is a pose on the trajectory corresponding to time $t$, and $p_{ij}(t)$ is the occupancy probability of cell $(i,j)$ as output by our spatiotemporal prediction network. This explicitly considers tracking by rewarding possible detections with uncertain positions over the planning horizon. The weighting factor $w$ determines the importance given to the two objectives, i.e., exploration and target tracking. In contrast to existing work [1, 25], our proposed target tracking utility promotes efficient target tracking by considering possible detections along the paths with the associated uncertain spatiotemporal variations in the target positions.
For planning, we use a sampling-based tree search to leverage these lookahead predictions. We consider a set of candidate paths $\Psi$ constructed using a finite set of heading changes, assuming a constant speed for the ASV. These paths are further parameterized with Bézier curves to generate smooth and feasible paths for the ASV. The planning objective as described in (9)-(10), is calculated for each of these candidate paths $\mathcal{T}$.
@@ V-A Experimental Setup
We perform simulations with an ASV in Gazebo with the VRX-simulator [30]. The wind is simulated as a stochastic process with the slowly varying component modeled as a first order Gauss-Markov process. We consider environments of size $100$ m × $\times\,100$ m discretized with size $\Delta_{x}=1$ m, $\Delta_{y}=1$ m. We generate spherical buoy markers in the simulations as targets of interest and model the wind forces on these targets as described in Fossen [31]. Different scenarios are generated by changing the mean wind speed, mean direction, and the spawn positions of the targets. Eight targets are used in all scenarios, and the mean wind speed is varied between 66-$10$ m/s. We run Monte Carlo simulations under these environmental conditions to perform a statistical study of the different components of our framework.
We consider the following metrics for comparison: the mean Shannon entropy in the occupancy map $H$, the mean squared error (MSE), and the mean detections ($\overline{N}$). The mean number of detections is computed as the average of the total number of detections, i.e., sum of the detections at each time step. The MSE between the ground truth and the occupancy grid [3] is computed as:
@@ V-B Ablation Study - Planning Utility
To study the effect of our spatiotemporal prediction network and the proposed planning utility, we compare various values of the weighting factor $w$ in (9), which determines the trade-off between the two objectives: reducing entropy in the occupancy grid and redetecting targets. Monte Carlo simulations are performed for w=0, 2, 5w=0,\,2,\,5 and 5(1t/B)5(1-t/B), and the resulting metrics are shown in Table I. Higher values of $w$ correspond to the planner prioritizing target redetection, while $w=0$ corresponds to an exploratory planner that only considers the entropy reduction in the occupancy grid. The ASV speed is set to $1.5$ m/s, and the planning horizon for the tree search is $25$ s. Each Monte Carlo trial is run for $B=250$ s.
w{w} H{H} $\overline{N}$} MSE
0 $\mathbf{0.661\pm 0.015}$ $1.123\pm 0.24$ $\mathbf{0.051\pm 0.008}$
22 $0.688\pm 0.020$ $1.529\pm 0.132$ $0.065\pm 0.012$
55 $0.703\pm 0.027$ $1.56\pm 0.326$ $0.073\pm 0.016$
5(1t/B)5(1-t/B) $0.681\pm 0.019$ $\mathbf{1.575\pm 0.24}$ $0.061\pm 0.007$
@@
Table I reveals that the purely exploratory case ($w=0$) is the best at reducing entropy and MSE. It is also worse at detecting targets by up to $\approx 29\%$ as compared to the other variants that use the information from the prediction network into the planning objective. This highlights the importance of our prediction network and the corresponding target tracking utility in the adaptive planning strategy to re-identify targets. We see that the trials with $w=5$ and w=5(1t/B)w=5(1-t/B) are nearly as effective at redetecting targets. These results demonstrate that our IPP utility improves target tracking performance by enabling more frequent target reacquisition.
@@ V-C Ablation Study - Mapping
Figure 5: Entropy and MSE for planning with the sampling-based planner, with and without the prediction step of the dynamic occupancy grid. The mapping accuracy is observed to be better by $21.8\%$ with the prediction step.
The entropy values at the end of the mission are similar, i.e., $3.82\%$ lower for the trials with the dynamic occupancy grid mapping with the prediction step, as compared to mapping without the prediction step. However, the MSE values are $21.8\%$ lower with our dynamic mapping approach as compared to the mapping without the prediction step. This indicates that the planner still effectively reduces uncertainty in the map in either case, as observed from the entropy. However, the mapping accuracy is better with the prediction step included, as the static occupancy grid does not account for the targets’ motion. This could also lead to multiple detections of a single target on the map. Such errors in the mapping also subsequently influence the decision-making of the robot, leading to suboptimal planning.
@@ V-D Comparison With Baselines
Figure 6: Entropy, MSE, and the mean detections $\overline{N}$ plotted with one standard deviation margin for the different planning strategies for the mission time ($250$ s). The proposed sampling-based planner is best at redetecting targets, closely followed by the receding horizon planner. Both these planning strategies benefit from longer lookahead predictions with the network. The greedy and sampling-based planners are both nearly as effective in reducing entropy.
We also evaluate our proposed sampling-based tree search planner by comparing it against other planning strategies. The factor $w$ is adaptively changed as 5(1t/B)5(1-t/B). We choose the following baselines:
  1. 1.
    An implementation of the tree search in a receding horizon manner, executing a path with only three out of the five waypoints that are used to parameterize the Bézier path before replanning.
  2. 2.
    Greedy planner that greedily considers only a single waypoint to calculate utility (9).
  3. 3.
    Coverage planning with lawnmower motion, that is traditionally used for monitoring [32].
  4. 4.
    A random planner that selects waypoints randomly.
The coverage and random planners are non-adaptive strategies that do not take new information into account during the decision-making process. The computational burden of the replanning step prohibits implementation of a receding horizon planner where only the first waypoint is traversed before planning again. The replanning steps in case of the receding horizon and the sampling-based planners take  $5$ s for the considered horizon of $25$ s and seven candidate paths, on a computer equipped with an i7-12700H processor and 32GB DDR5 RAM.
The mean entropy $H$, mean detections $\overline{N}$, and the MSE values for the various planning strategies are plotted in Fig. 6. We observe that the non-adaptive baselines perform the worst in exploring the map and also at redetecting targets. The greedy planner and the sampling-based planner report very similar scores in terms of entropy and MSE measures, indicating they are equally effective at exploring the environment. However, the sampling-based planner is better at tracking targets, as it detects $16.8\%$ more targets than the greedy planner, and $4.3\%$ more targets as compared to the receding horizon planner at the end of mission time. This can be attributed to the sampling-based and receding horizon planners computing the utility by predicting sensor measurements along the path for the entire planning horizon, as opposed to the greedy planner, which greedily considers only the final waypoint on the path. From the entropy plot, we see that the receding horizon planner is not as effective in exploring the map as compared to the other adaptive strategies, due to the additional computational burden due to more frequent replanning, as the ASV has to stop and replan more frequently in this case.
@@ VI Field Tests
To validate our framework in a real-world scenario, we perform field tests with an ASV for mapping freely drifting spherical floating buoy markers, as shown in Fig. 1. The ASV is overactuated with four thrusters in the X-configuration. It is equipped with a Cube Orange IMU and a Here4 GNSS for navigation. It has a Jetson Orin Nano, and software is implemented with ROS2 as middleware. Trials are performed on a $25$ m× $\times\,25$ m area for a mission time of $150$ s. We use a ZED 2i camera with $72^{\circ}$ horizontal fov. We consider detections up to a distance of $15$ m, as the camera gives reliable depth estimates up to 1515-2020 m. We use similar inverse sensor models for mapping as described in Sec. IV-B, with the parameters scaled accordingly. The ASV’s advance speed is set to $0.5$ m/s, and the lookahead time for the planning step is $25$ s.
Figure 7: Evolution of the mean entropy $H$ and mean detections $\overline{N}$ during our field trial missions. The trial with $w=5$ is most effective at redetecting targets.
During the experiments, moderate winds of $4$ m/s were observed from West to East. We place four buoy markers near the left edge of the map at the start of each trial, and they are found to drift freely with the wind towards East due to the wind. We perform three trials with the sampling-based planner, each corresponding to $w=0,5$ and 5(1t/B)5(1-t/B), respectively, and discuss their results here. The resulting entropy and mean detections from the field trials are plotted in Fig. 7. It is observed that the trial with $w=5$ is most effective at redetecting targets, as evident from several peaks in the graph for $\overline{N}$. These results align with simulations showing that higher values of $w$ promote target tracking, further highlighting our proposed utility. We also show the occupancy maps at various instants for this trial in Fig. 8. The ASV first detects three of the four targets at $t=15$ s. We notice that at later intervals, $t=90$ s and $145$ s, the ASV tracks all four targets. An accompanying video of this experiment is also provided [^1] Video from the field trials with the ASV -https://imgur.com/a/O9aycaA.

### v3

@@ IV-D Spatiotemporal Prediction Network
Maintaining an accurate map of target positions requires planning informative trajectories for the robot that facilitate repeated detections of targets. This involves predicting changes in the map over some time frame. We introduce a spatiotemporal prediction network to predict the uncertainty of the target positions in the map over a certain time horizon. Although our dynamic occupancy mapping method contains a prediction step to estimate the drift of targets, it is not accurate at predicting drift for longer time horizons, due to the stochastic nature of environmental disturbances. Therefore, we create a formulation to estimate the spatial and temporal variations of the target positions for future time steps.
We assume that the position of each target is represented by a Gaussian kernel, with the uncertainty in its position estimate growing with time, similar to previous work [wang2023spatio]. To reflect this assumption, we assume that the variances of the kernel are proportional to the wind speed and increase with the prediction time, with the principal axis being along the direction of the measured wind. We assume that the centers of each kernel translate by distances calculated as per (6). The variance along the direction of the wind is assumed to be greater, as for most realistic scenarios, the wind speed varies faster than the direction [kaimal1994atmospheric]. Based on this assumption, the standard deviations of the kernel along the major and minor axes, $\sigma_{\parallel}$, and $\sigma_{\perp}$ for a prediction interval $t$ are set as $\sigma_{\parallel}=0.5\,\gamma v_{w}t$, and $\sigma_{\perp}=0.2\,\gamma v_{w}t$ respectively. In case of negligible wind speeds ($v_{w}\approx 0$), $\sigma_{\parallel}=\sigma_{\perp}=0.1\,t$.
@@
We provide the wind parameters $v_{w}$, $\psi_{w}$, and the prediction interval as other inputs to the network. The network output is a grid of the same size with predicted position uncertainties of the targets at an arbitrary time step. We use a UNet architecture [ronneberger2015u], which is widely used for image-to-image prediction tasks such as segmentation. The wind parameters and the prediction time are embedded into the latent space of the network.
To train our network, a synthetic dataset is created to simulate input conditions with different wind conditions, target positions, and prediction intervals. To construct inputs resembling target occupancies in the map, binary grids of size $100\times 100$ containing an arbitrary number of objects (<20<20) with random positions are generated. For each of these input grids representing the spatial distribution of the targets, outputs are computed as described above, for various combinations of wind parameters and prediction time steps. While predicting the drift of objects on the water surface is a difficult problem [breivik2011wind], we find that these assumptions suit our problem setup. Using an input representation that is agnostic to the number of targets makes our method generalizable, as validated through both simulation and field experiments, in Secs. V and VI.
@@ IV-E Informative Planning for Dynamic Target Tracking
The robot needs to replan at frequent intervals to account for the changing environment. It must explore unvisited regions of the map while redetecting targets. Planning informative trajectories in dynamic environments necessitates predicting both spatial and temporal variations in the environment. Commonly used IPP objectives are based on Shannon entropy measured based on the map state [popovic2020informative, moon2022tigris]. However, these formulations do not explicitly account for temporal variations and periodic target redetections. In contrast, we exploit the predictions from our spatiotemporal network to include the temporal uncertainties in target positions in the planning objective.
To trade off between exploring the environment and redetecting targets, we introduce a new planning utility that consists of both exploration and target tracking objectives. We propose a new target tracking component in the utility that explicitly considers possible target detections by estimating the spatiotemporal variations in the targets’ position estimates in the planning objective. It is computed with the predictions from our spatiotemporal network, for prediction steps over the considered planning horizon. To compute the predicted change in entropy, the dynamic occupancy grid mapping is forward-simulated for predicted measurements along considered trajectories. We define the planning utility as:
$I(\mathcal{T})$ $=\Big[H(\mathcal{M})-H(\mathcal{M}\,|\,z_{\mathcal{T}})\Big]+\frac{w}{|\mathcal{T}|}\sum_{t}{J(\boldsymbol{x}^{t})}\,$, (9)
@@
where $H(\mathcal{M})$ denotes the mean Shannon entropy of the occupancy grid, and $H(\mathcal{M}|z_{\mathcal{T}})$ denotes the predicted posterior mean entropy after simulating measurements $z_{\mathcal{T}}$ along a trajectory $\mathcal{T}=\{\boldsymbol{x}^{t}|t=1,\dots,|\mathcal{T}|\}$. The target tracking objective is defined as the averaged sum of the utility $J(\boldsymbol{x}^{t})$ over the prediction steps, where $\boldsymbol{x}^{t}$ is a pose on the trajectory corresponding to time $t$, and $p_{ij}(t)$ is the occupancy probability of cell $(i,j)$ as output by the network. This explicitly considers tracking by rewarding possible detections with uncertain positions over the planning horizon. The weighting factor $w$ determines the importance given to the two objectives, i.e., exploration and target tracking. In contrast to existing work [meera2019obstacle, moon2022tigris], our proposed target tracking utility promotes efficient target tracking by considering possible detections along the paths with the associated uncertain spatiotemporal variations in the target positions.
For planning, we use a sampling-based planner to leverage these lookahead predictions. We consider a set of candidate paths $\Psi$ constructed using a finite set of evenly discretized heading changes, assuming a constant speed for the ASV. These paths are smoothed with Dubins curves to generate dynamically feasible paths for the ASV. The planning objective as described in (9)-(10), is calculated for each of these candidate paths $\mathcal{T}$.
@@ V-A Experimental Setup
We perform simulations with an ASV in Gazebo with the VRX simulator [bingham2019toward]. The wind is simulated as a stochastic process with the slowly-varying component modeled as a first-order Gauss-Markov process. We consider environments of size $100$ m × $\times\,100$ m discretized with size $\Delta_{x}=1$ m, $\Delta_{y}=1$ m. We generate spherical buoy markers in the simulations as targets of interest and model the wind forces on these targets as described by Fossen [fossen2011handbook]. Different scenarios are generated by changing the mean wind speed, mean direction, and the spawn positions of the targets. Eight targets are present in the environment, and the mean wind speed is varied between 66-$10$ m/s in the scenarios. We run Monte Carlo simulations under these environmental conditions to perform a statistical study of the different components of our framework. The mapping is performed at $5$ Hz. The planning horizon for replanning is $25$ s, with time steps of $1$ s. The replanning step for seven candidate paths takes $\approx 0.08$ s, on a computer equipped with an i7-12700H processor, 32GB DDR5 RAM and RTX3050 GPU, for parallelized computations of the utility (9). Each Monte Carlo trial is run for a mission time of $B=250$ s, and the ASV’s speed is set to $1.5$ m/s. For path-following control, line of sight guidance [fossen2011handbook] is used along with PID speed and heading controllers.
We consider the following metrics for comparison: the reduction in the mean Shannon entropy in the occupancy map ($\Delta H$), the mean squared error (MSE), and the mean detections ($\overline{N}$). The mean number of detections is computed as the average of the total number of detections, i.e., sum of the detections at each time step. The MSE between the ground truth and the occupancy grid is computed as [barrionuevo2025optimizing]:
@@ V-B Ablation Study - Planning Utility
To study the effect of our spatiotemporal prediction network and the proposed planning utility, we compare various values of the weighting factor $w$ in (9), which determines the trade-off between the two objectives: reducing entropy in the occupancy grid and redetecting targets. Monte Carlo simulations are performed for w=0, 2, 5, 5t/Bw=0,\,2,\,5,\,5\,t/B, and 5(1t/B)5(1-t/B), with the sampling-based planner executed in a receding horizon fashion. The resulting metrics are shown in Table I. Higher values of $w$ correspond to the planner prioritizing target redetection, while $w=0$ corresponds to an exploratory planner that only considers the entropy reduction in the occupancy grid in the utility.
w{w} $\Delta H$} $\overline{N}$} MSE
0 $\mathbf{0.345\pm 0.012}$ $1.122\pm 0.307$ $\mathbf{0.046\pm 0.005}$
22 $0.314\pm 0.019$ $1.594\pm 0.193$ $0.065\pm 0.011$
55 $0.295\pm 0.019$ $1.690\pm 0.142$ $0.076\pm 0.011$
5t/B5\,t/B $0.287\pm 0.014$ $1.670\pm 0.131$ $0.081\pm 0.008$
5(1t/B)5(1-t/B) $0.293\pm 0.021$ $\mathbf{1.696\pm 0.137}$ $0.077\pm 0.012$
@@
Table I reveals that the purely exploratory case ($w=0$) is the best at reducing entropy and subsequently MSE. It is also worse at detecting targets by at least $\approx 31\%$ as compared to the other variants that use the tracking information gain in the planning objective. This highlights the importance of our prediction network and the corresponding target tracking utility in the adaptive planning strategy to re-identify targets. Further, we observe that higher values of $w$ correspond to more target detections. These results demonstrate that our IPP utility improves target tracking performance by enabling more frequent target reacquisition.
@@ V-C Ablation Study - Mapping
Figure 5: Entropy reduction $\Delta H$ and MSE for planning with the sampling-based planner, with and without the prediction step of the dynamic occupancy grid. With the prediction step, the mapping accuracy is observed to be better by $21.8\%$.
The reduction in entropy at the end of the mission is similar, i.e., the trials including the prediction step in the dynamic occupancy grid mapping $8.4\%$ , as compared to mapping without the prediction step. However, the MSE values are $21.8\%$ lower with our dynamic mapping approach as compared to the mapping without the prediction step. This indicates that the planner still effectively reduces uncertainty in the map in either case, as observed from the entropy. However, the mapping accuracy is better with the prediction step included, as the static occupancy grid does not account for the targets’ motion. This could also lead to multiple detections of a single target on the map. Such errors in the mapping also subsequently influence the decision-making of the robot, leading to suboptimal planning.
@@ V-D Comparison With Baselines
N = 8 N = 16 N = 32
$\Delta H$ $\overline{N}$ $\Delta H$ $\overline{N}$ $\Delta H$ $\overline{N}$
Sampling-based $0.315\pm 0.014$ $1.469\pm 0.296$ $0.304\pm 0.016$ $2.539\pm 0.329$ $\mathbf{0.305\pm 0.023}$ $5.158\pm 1.016$
Receding horizon $0.295\pm 0.019$ $\mathbf{1.690\pm 0.142}$ $0.266\pm 0.014$ $\mathbf{3.015\pm 0.268}$ $0.258\pm 0.012$ $\mathbf{6.089\pm 1.014}$
Greedy $\mathbf{0.346\pm 0.013}$ $1.335\pm 0.247$ $\mathbf{0.315\pm 0.026}$ $2.490\pm 0.416$ $0.302\pm 0.017$ $4.244\pm 0.651$
Coverage $0.239\pm 0.005$ $0.555\pm 0.43$ $0.242\pm 0.006$ $1.389\pm 0.904$ $0.236\pm 0.003$ $1.809\pm 0.159$
Random $0.179\pm 0.064$ $0.706\pm 0.49$ $0.171\pm 0.059$ $1.407\pm 0.93$ $0.136\pm 0.043$ $2.383\pm 1.460$
TABLE II: Entropy, and mean detections $\overline{N}$ compared for the considered planning strategies for cases with various numbers of targets. The receding horizon planner is best at redetecting targets, followed by the sampling-based planner. Both these planning strategies benefit from longer lookahead predictions with the network. The greedy and sampling-based planners are both nearly as effective in reducing entropy.
We also evaluate our proposed sampling-based planner by comparing it against other planning strategies. It is executed in both a finite horizon and a receding horizon manner. In the receding horizon implementation, replanning is done after $50\%$ of the previous path is traversed. We choose the following baselines: (i) a greedy planner that greedily considers only a single waypoint to calculate utility (9); (ii) coverage planning with lawnmower motion, traditionally used for monitoring [galceran2013survey]; (iii) a random planner that selects waypoints randomly. The coverage and random planners are non-adaptive strategies that do not take new information into account during the decision-making process. For the other planning strategies, the factor $w$ is set to $5$. Experiments are also performed for scenarios with a varying number of objects $N=8,16,32$ to evaluate the ability of our approach to generalize to larger numbers of targets.
The mean entropy $H$ and mean detections $\overline{N}$ for the various planning strategies are shown in Table II. We observe that the non-adaptive baselines perform the worst in exploring the map and also in redetecting targets. The greedy planner reports very similar scores in terms of entropy as compared to the sampling-based planner, indicating it is equally effective at exploring the environment. However, the receding horizon planner is the best at tracking targets in all cases, followed by the sampling-based planner. This can be attributed to the sampling-based and receding horizon planners computing the utility by predicting sensor measurements along the path for the entire planning horizon, as opposed to the greedy planner that considers only a single waypoint rather than optimizing the utility over a path.
@@ VI Field Tests
To validate our framework in a real-world scenario, we perform field tests with an ASV for mapping freely drifting spherical buoy markers, as shown in Fig. 1. The ASV is overactuated with four thrusters in the X-configuration. It is equipped with a Cube Orange IMU and a Here4 GNSS for navigation. It has a Jetson Orin Nano, and software is implemented in ROS2. Trials are performed on a $25$ m× $\times\,25$ m area for a mission time of $150$ s. We use a ZED 2i camera with $72^{\circ}$ horizontal FOV. Detections up to a distance of $15$ m are considered, as the camera gives reliable depth estimates up to 1515-2020 m. We use similar inverse sensor models for mapping as described in Sec. IV-B, with parameters scaled accordingly. The ASV’s advance speed is set to $0.5$ m/s, and the lookahead time for the planning is $25$ s.
Figure 6: Evolution of $\Delta H$ and mean detections $\overline{N}$ during field trials, where $w=5$ is most effective at redetecting targets.
During the experiments, moderate winds of $4$ m/s were observed from West to East. We place four buoy markers near the left edge of the map at the start of each trial, and they are found to drift freely with the wind towards East due to the wind. We perform three trials with the sampling-based planner, each corresponding to $w=0,5$, and 5(1t/B)5(1-t/B), respectively, and discuss their results here. The resulting entropy and mean detections from the field trials are plotted in Fig. 6. It is observed that the trial with $w=5$ is most effective at redetecting targets, as evident from several peaks in the graph for $\overline{N}$. These results align with simulations showing that higher values of $w$ promote target reacquisition, further highlighting the usefulness of our proposed utility. We also show the occupancy maps at various instants for this trial in Fig. 7. The ASV first detects three of the four targets at $t=15$ s. We notice that at later intervals, $t=90$ s and $145$ s, the ASV tracks all four targets. An accompanying video of this experiment is also provided [^1] A video from the field trials with the ASV is available at https://youtu.be/nT_s3GbHHKU..